On the first day of school, the percentage of boys in a particular class is 60%. During the school year, six girls move away, and are replaced in the class by six boys; this makes the class roster 75% boys. Find the number of boys and girls in the class on the first day of school.
Answer by checkley71(8403)

b=.6t b=boys t=total number in class
b+6=.75t substituting .6t for b in the second equation we get
.6t+6=.75t ot 6=.15t or t=6/.15 or t=40 therefore .6*40=boys or .6*40=24 boys and 16 girls (40-24) at the beginning of school.
24+6=30 OR .75*40=30